OpenStudy (anonymous):

5g + 4(-5 + 3g) = 1 - g

OpenStudy (hba):

\[5g-20+12g=1-g\]

OpenStudy (hba):

\[17g-20=1-g\]

OpenStudy (hba):

\[18g=21\]

OpenStudy (hba):

\[g=\frac{ 21 }{ 18 }=\frac{ 7 }{ 6 }\]

OpenStudy (hba):

@tarachery

OpenStudy (anonymous):

thank youu =)